# How to Find the Exact Pareto Front for Multi-Objective MDPs?

### v1

@@ Appendix F Locality Property of boundary of ${\mathbb{J}}(\mu)$ and Pareto front
Let ${\mathbb{F}}_{P}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1,\text{ND}}}\cup\{J^{\pi}\}))$ denote the set of faces of ${\mathbb{F}}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1,\text{ND}}}\cup\{J^{\pi}\}))$ that satisfying Lemma 3. Then, we have
@@ Appendix G Find Pareto front on CCS
Given a convex polytope ${\mathbb{C}}$, a face ${\mathbb{F}}$ of ${\mathbb{C}}$ lies on the Pareto front if the following conditions hold:
  • If ${\mathbb{F}}$ is a facet of ${\mathbb{C}}$, i.e., $\text{dim}({\mathbb{F}})=\text{dim}({\mathbb{C}})-1$, and can be written as ${\mathbb{F}}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}^{\top}{\bm{x}}=c_{0}\}$, then ${\mathbb{F}}$ lies on the Pareto front if ${\bm{w}}>{\bm{0}}$.
  • If $10\leq\text{dim}({\mathbb{F}})<\text{dim}({\mathbb{C}})-$ 1, and ${\mathbb{F}}$ is the intersection of facets of ${\mathbb{C}}$, i.e., ${\mathbb{F}}=\cap_{i}{\mathbb{F}}_{i}$, where each facet is written as ${\mathbb{F}}_{i}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}_{i}^{\top}{\bm{x}}=c_{i}\}$, then ${\mathbb{F}}$ lies on the Pareto front if there exists a linear combination of the normal vectors with weight vector $\boldsymbol{\alpha}\geq 0$, such that $\sum_{i}\boldsymbol{\alpha}_{i}{\bm{w}}_{i}>0$.
@@
When ${\mathbb{F}}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}^{\top}{\bm{x}}=c_{0}\}$ is a facet of ${\mathbb{C}}$, by the definition of a valid linear inequality, the inequality ${\bm{w}}^{\top}{\bm{x}}\leq c_{0}$ holds for all points ${\bm{x}}\in{\mathbb{C}}$. Since ${\bm{w}}>0$, for any ${\bm{x}}$ on ${\mathbb{F}}$ and any ${\bm{y}}\in{\mathbb{C}}$, we have ${\bm{w}}^{\top}({\bm{y}}-{\bm{x}})\leq 0$. By Lemma 8, this implies that no point in ${\mathbb{C}}$ dominates ${\bm{x}}$, so ${\mathbb{F}}$ lies on the Pareto front.
@@
where $\boldsymbol{\alpha}\geq 0$. Since $\sum_{i}\boldsymbol{\alpha}_{i}{\bm{w}}_{i}>0$, for any ${\bm{x}}$ on ${\mathbb{F}}$ and any ${\bm{y}}\in{\mathbb{C}}$, there exists a vector ${\bm{w}}>0$ such that ${\bm{w}}^{\top}({\bm{y}}-{\bm{x}})\leq 0$. Therefore, by Lemma 8, no point in ${\mathbb{C}}$ strictly dominates ${\bm{x}}$, and ${\mathbb{F}}$ lies on the Pareto front. ∎
@@
Therefore, $\Pi_{F}=\arg\max_{\pi\in\Pi_{D}}{\bm{w}}^{\top}J^{\pi}$, where $\Pi_{D}$ denotes the set of all deterministic policies. By Lemma 13, the convex hull of $\Pi_{D}$, i.e., $\left\{\pi\mid\pi(a|s)=\sum_{\pi_{i}\in\Pi_{F}}\boldsymbol{\alpha}_{i}\pi_{i}(a|s),\boldsymbol{\alpha}\geq 0,\sum_{i}\boldsymbol{\alpha}_{i}=1\right\}$, constructs all policies (including stochastic and deterministic) that maximize ${\bm{w}}^{\top}J^{\pi}$. Thus, the long-term returns of $\left\{\pi\mid\pi(a|s)=\sum_{\pi_{i}\in\Pi_{F}}\boldsymbol{\alpha}_{i}\pi_{i}(a|s),\boldsymbol{\alpha}\geq 0,\sum_{i}\boldsymbol{\alpha}_{i}=1\right\}$ construct $F$. Formally, we have
$F=\left\{J^{\pi}\mid\pi(a|s)=\sum_{\pi_{i}\in\Pi_{F}}\boldsymbol{\alpha}_{i}\pi_{i}(a|s),\boldsymbol{\alpha}\geq 0,\sum_{i}\boldsymbol{\alpha}_{i}=1\right\}.$
This concludes the proof. ∎

### v2

@@ Appendix F Locality Property of boundary of ${\mathbb{J}}(\mu)$ and Pareto front
(Restatement of Proposition 1) Let ${\mathbb{F}}_{P}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1,\text{ND}}}\cup\{J^{\pi}\}))$ denote the set of faces of ${\mathbb{F}}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1,\text{ND}}}\cup\{J^{\pi}\}))$ that satisfying Lemma 3. Then, we have
@@ Appendix G Find Pareto front on CCS
(Restatement of Lemma 3) Given a convex polytope ${\mathbb{C}}$, a face ${\mathbb{F}}$ of ${\mathbb{C}}$ lies on the Pareto front if and only if the following conditions hold:
  • When ${\mathbb{F}}$ is a facet of ${\mathbb{C}}$, i.e., $\text{dim}({\mathbb{F}})=\text{dim}({\mathbb{C}})-1$, there exists ${\bm{w}}>{\bm{0}}$ such that ${\mathbb{F}}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}^{\top}{\bm{x}}=c_{0}\}$.
  • When $10\leq\text{dim}({\mathbb{F}})<\text{dim}({\mathbb{C}})-$ 1, and ${\mathbb{F}}$ is the intersection of facets of ${\mathbb{C}}$, i.e., ${\mathbb{F}}=\cap_{i}{\mathbb{F}}_{i}$, where each facet is written as ${\mathbb{F}}_{i}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}_{i}^{\top}{\bm{x}}=c_{i}\}$, there exists a linear combination of the normal vectors with weight vector $\boldsymbol{\alpha}\geq 0$, such that $\sum_{i}\boldsymbol{\alpha}_{i}{\bm{w}}_{i}>0$.
@@
When ${\mathbb{F}}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}^{\top}{\bm{x}}=c_{0}\}$ is a facet of ${\mathbb{C}}$, by the definition of a valid linear inequality, the inequality ${\bm{w}}^{\top}{\bm{y}}\leq c_{0}$ holds for all points ${\bm{y}}\in{\mathbb{C}}$. Since ${\bm{w}}>0$, for any ${\bm{x}}$ on ${\mathbb{F}}$ and any ${\bm{y}}\in{\mathbb{C}}$, we have ${\bm{w}}^{\top}({\bm{y}}-{\bm{x}})\leq 0$. By Lemma 8, this is equivalent to that no point in ${\mathbb{C}}$ dominates ${\bm{x}}$ and ${\mathbb{F}}$ lies on the Pareto front.
@@
where $\boldsymbol{\alpha}\geq 0$. Since $\sum_{i}\boldsymbol{\alpha}_{i}{\bm{w}}_{i}>0$, for any ${\bm{x}}$ on ${\mathbb{F}}$ and any ${\bm{y}}\in{\mathbb{C}}$, there exists a vector ${\bm{w}}>0$ such that ${\bm{w}}^{\top}({\bm{y}}-{\bm{x}})\leq 0$. Therefore, by Lemma 8, this is equivalent to no point in ${\mathbb{C}}$ strictly dominates ${\bm{x}}$ and ${\mathbb{F}}$ lies on the Pareto front.
@@
Therefore, $\Pi_{F}=\arg\max_{\pi\in\Pi_{D}}{\bm{w}}^{\top}J^{\pi}$, where $\Pi_{D}$ denotes the set of all deterministic policies. By Lemma 13, the convex hull of $\Pi_{D}$, i.e., $\left\{\pi\mid\pi(a|s)=\sum_{\pi_{i}\in\Pi_{F}}\boldsymbol{\alpha}_{i}\pi_{i}(a|s),\boldsymbol{\alpha}\geq 0,\sum_{i}\boldsymbol{\alpha}_{i}=1\right\}$, constructs all policies (including stochastic and deterministic) that maximize ${\bm{w}}^{\top}J^{\pi}$. Thus, the long-term returns of $\left\{\pi\mid\pi(a|s)=\sum_{\pi_{i}\in\Pi_{F}}\boldsymbol{\alpha}_{i}\pi_{i}(a|s),\boldsymbol{\alpha}\geq 0,\sum_{i}\boldsymbol{\alpha}_{i}=1\right\}$ construct $F$. Formally, $F=\left\{J^{\pi}\mid\pi(a|s)=\sum_{\pi_{i}\in\Pi_{F}}\boldsymbol{\alpha}_{i}\pi_{i}(a|s),\boldsymbol{\alpha}\geq 0,\sum_{i}\boldsymbol{\alpha}_{i}=1\right\}$. This concludes the proof. ∎
Appendix H Complexity Comparison between OLS and Algorithm 1
Let $S$, $A$, and $D$ represent the state space size, the action space size, and the reward dimension, respectively. Let $N$ denote the number of vertices on the Pareto front.
For each iteration, OLS solves the single-objective MDP with a preference vector, removes obsolete preference vectors with the new points, and computes new preference vectors. Solving the single-objective planning problem incurs a complexity of $\mathcal{O}(C_{\text{so}}+C_{\text{pe}})$, where $C_{\text{so}}$ and $C_{\text{pe}}$ represent the computational complexity of single-objective MDP solver and policy evaluation, respectively. The complexity of comparing with the previous candidate Pareto optimal policies and removing obsolete ones is $\mathcal{O}(ND)$, and the complexity of calculating a new preference vector is denoted as $C_{\text{we}}$. The computational complexity of OLS is $\mathcal{O}(N^{2}D+NC_{\text{we}}+NC_{\text{so}}+NC_{\text{pe}})$.
Algorithm 1 requires only a one-time single-objective planning step during the initialization phase to obtain the initial deterministic Pareto-optimal policy. The overall computational complexity of our algorithm is $\mathcal{O}(C_{\text{so}}+NSAC_{\text{pe}}+NC_{\text{ND}}+NC_{\text{cvx}})$, where $C_{\text{ND}}$ and $C_{\text{cvx}}$ represent the computational complexity to select the non-dominated policies and computing faces of the convex hull, respectively.
Specifically, we dive deeper into the computational complexity of each module.
  • $C_{\text{so}}$. The computational complexity of single-objective planning is $C_{\text{so}}=\mathcal{O}\left(S^{4}A+S^{3}A^{2}\right)$.
  • $C_{\text{pe}}$. The computational complexity of the policy evaluation is $C_{\text{pe}}=\mathcal{O}(S^{3}D)$.
  • $C_{\text{ND}}$. $C_{\text{ND}}$ refers to the complexity of finding $\Pi_{1,\text{ND}}$, the set of non-dominated policies from $\Pi_{1}$. Let $N_{\text{ND}}=|\Pi_{1,\text{ND}}|=$. The non-dominated policies searching complexity is $C_{\text{ND}}=\mathcal{O}(SAN_{\text{ND}}D)$.
  • $C_{\text{cvx}}$. $C_{\text{cvx}}$ is the complexity of computing the convex hull faces formed by $N_{\text{ND}}+1$ points that intersect at the current policy. If the number of vertices of the output convex hull is $N_{\text{cvx}}$. By [1], the complexity of the calculation of convex hull is $\mathcal{O}(N_{\text{ND}}\log N_{\text{cvx}})$ when $D\leq 3$ and $\mathcal{O}(N_{\text{ND}}f_{N_{\text{cvx}}}/N_{\text{cvx}})$, where $f_{L}=\mathcal{O}(L^{\lfloor D/2\rfloor}/(\lfloor D/2\rfloor!))$.
As Algorithm 1 directly searches the Pareto front, the total number of iterations corresponds to the number of vertices on the Pareto front, avoiding the quadratic complexity introduced by the preference vector updates in OLS. Moreover, note that calculating new weight in OLS requires calculating the vertices of the convex hull constructed by $N+D+2$ hyperplanes of dimension $D+1$, making it significantly more expensive than the local convex hull construction of neighboring policies in our proposed algorithm. Hence, $C_{we}$ is much larger than $C_{\text{cvx}}$.
In conclusion, our algorithm induces total computational complexity of $\mathcal{O}(S^{4}A+S^{3}A^{2}+NS^{4}A+NSAN_{\text{ND}}D+C_{\text{cvx}})$, while OLS has computation complexity $\mathcal{O}(N(S^{4}A+S^{3}A^{2})+NS^{3}+N^{2}D+NC_{\text{we}})$. Given $C_{\text{cvx}}<C_{\text{we}}$, our algorithm is more efficient than OLS in all cases.
Appendix I Supporting Algorithms
For completeness, we provide the PPrune algorithm [13], which identifies the non-dominated vector value functions ${\mathbb{V}}^{*}$ from a given set of vector value functions ${\mathbb{V}}$. Note that, with slight abuse of notation, ${\mathbb{V}}$ here refers to the set of vector value functions rather than the set of vertices used in the main text.
Algorithm 3 PPrune(${\mathbb{V}}$[13]
1:Input: A set of value vectors ${\mathbb{V}}$
2:${\mathbb{V}}^{*}\leftarrow\varnothing$
3:while ${\mathbb{V}}\neq\varnothing$ do
4:     $V\leftarrow$ the first element of ${\mathbb{V}}$
5:     for all $V^{\prime}\in{\mathbb{V}}$ do
6:         if $V^{\prime}\succ V$ then
7:              $V\leftarrow V^{\prime}$
8:         end if
9:     end for
10:     Remove $V$ and all vectors dominated by $V$ from ${\mathbb{V}}$
11:     Add $V$ to ${\mathbb{V}}^{*}$
12:end while
13:return ${\mathbb{V}}^{*}$
The benchmark algorithm retrieves the Pareto front in two steps: (1) evaluate all deterministic policies and compute the convex hull of all non-dominated deterministic policies and (2) apply Algorithm 2 to identify the Pareto front. The details of the benchmark algorithm are shown in Algorithm 4.
Algorithm 4 Benchmark Algorithm
3:Generate all deterministic policies $\Pi_{D}$.
4:Select non-dominated policies $\Pi_{D,\text{ND}}$ from $\Pi_{D}$.
5:Calculate the convex hull formed by $\{J^{\pi}\mid\pi\in\Pi_{D,\text{ND}}\}$.
6:Extract Pareto optimal faces $\mathbb{P}(\Pi)$ and vertices ${\mathbb{V}}(\Pi)$ from the convex hull with Algorithm 2.
Appendix J Experiment Results
We adopt the deep-sea-treasure setting from [10] and construct a scenario where the agent navigates a grid world, making trade-offs between different objectives. At each step, the agent can move up, down, left, or right, and the reward for each grid is generated randomly. We compare the performance of the OLS algorithm and the proposed Pareto front searching algorithm under different grid sizes (varying numbers of rows and columns). The experiments were conducted on smaller grid sizes to ensure that the running time of OLS remained manageable. As shown in the Fig. 8, the proposed algorithm consistently outperforms OLS in terms of efficiency across all tested settings.
We evaluate the trend of the number of vertices $N$ in terms of $S$ and $A$ with the state space size ranging from 10 to 50 and the action space sizes ranging from 10 to 20. The reward dimension is 3. As shown in Fig. 8, the result shows that the number of vertices grows approximately as $N\approx kS^{3}A$, where $k$ is a constant number. Combined with the computational complexity analysis $\mathcal{O}(S^{4}A+S^{3}A^{2}+NS^{4}A+NSAN_{\text{ND}}D+C_{\text{cvx}})$, this demonstrates the scalability of our algorithm.
Figure 7: Pareto front of a simple MDP with $S=4$, $A=3$, and $D=3$.